OpenStudy (anonymous):

How do I solve this equation: lnx=ln3-ln(x+5)?

OpenStudy (anonymous):

start with \[\ln(x+5)+\ln(x)=\ln(3)\]

OpenStudy (anonymous):

ohh i see it, thank you!